# Broken intrinsic symmetry induced magnon-magnon coupling in synthetic ferrimagnets

### v1

@@ III Results and Discussion
The relationship between the saturation magnetizations of the ferromagnetic layers can be established using the normalized magnetization of the forced ferromagnetic ($M_{\mathrm{ff}}$) and fully antiferromagnetic ($M_{\mathrm{afm}}$) states [magnetization at point ‘2’ to point ‘3’ in Fig. 1(c), respectively]. The ratio of the saturation magnetizations can be expressed as follows:
$\mathrm{\frac{M_{S2}}{M_{S1}}=\frac{M_{ff}-M_{afm}}{M_{ff}+M_{afm}}}.$ (4)
For a Ru thickness of $0.8$ nm, we find $M_{\mathrm{ff}}=1.0$ and $M_{\mathrm{afm}}=0.29$. The saturation magnetization used in this work for NiFe was taken from literature to be 10.1 kOe Mamin et al. (1989), which allows us to calculate the saturation magnetization of CoFe to be 18.5 kOe using Eq. (4). From magnetometry measurements in Fig. 1(c) we found $H_{\mathrm{0,ff}}$ and $H_{\mathrm{0,afm}}$ to be 590 Oe and 155 Oe, respectively. This provides values for $J_{\mathrm{q}}$ and $J_{\mathrm{bq}}$ to be $2.95\leavevmode\nobreak\ \mathrm{kJ/m^{3}}$ and $0.74\leavevmode\nobreak\ \mathrm{kJ/m^{3}}$, respectively, using Eqs. (3a) and (3b).
Figure 4: (a) Simulated coupling gap $g$ (blue) and simulated degeneracy field $H_{\mathrm{g}}$ (green) vs. quadratic interaction constant $J_{\mathrm{q}}$ at a fixed biquadratic interaction constant $J_{\mathrm{bq}}=0.74\leavevmode\nobreak\ \mathrm{kJ/m^{3}}$. The degeneracy field $H_{\mathrm{g}}$ was determined as the field where the high- and low-frequency modes have the smallest separation in frequency ($g$).Solid lines represent linear fits with respect to $J_{\mathrm{q}}$.(b) Simulated coupling gap $g$ (blue) and simulated degeneracy field $H_{\mathrm{g}}$ (green) vs. biquadratic interaction constant $J_{\mathrm{bq}}$ at a fixed quadratic interaction constant $J_{\mathrm{q}}=29.5\leavevmode\nobreak\ \mathrm{kJ/m^{3}}$. Solid lines represent linear fits with respect to $J_{\mathrm{bq}}$. Red points indicate the simulation conditions resembling the experimental conditions for a Ru thickness of 0.8 nm. (c) Experimentally observed gap vs. degeneracy field (green) and simulated gap vs. degeneracy field (blue). The experimental data points are extracted from the BLS measurements. STFMR measurements show a similar linear trend for gap vs. degeneracy field.
To gain a better understanding of the experimental results, we performed micromagnetic modeling using the graphics processor unit (GPU)-accelerated program Mumax3 Vansteenkiste et al. (2014); Exl et al. (2014). Our structure was modeled as $128\times 64\times 3$ cells with an individual cell size of $4.0$ nm × $\times\leavevmode\nobreak\ 4.0$ nm × $\times\leavevmode\nobreak\ 4.0$ nm, with the middle layer representing the spacer. Each layer was defined into a region to set material parameters separately. The values for the exchange stiffness of CoFe and NiFe used in the simulation were $3.0\times 10^{-11}$ J/m and $1.3\times 10^{-11}$ J/m, respectively Parreiras and Martins (2015). This corresponds to exchange lengths for NiFe and CoFe of 5.29 nm and 4.32 nm, respectively Parreiras and Martins (2015). Crystalline anisotropy was added for the CoFe layer with a crystalline anisotropy constant of $K=-30\times 10^{3}$ J/m3 along the (1 0 0) direction Parreiras and Martins (2015). The following Gilbert damping constants were used for CoFe and NiFe, respectively: $\mathrm{\alpha_{\mathrm{CoFe}}=0.005}$ Weber et al. (2019) and $\mathrm{\alpha_{\mathrm{NiFe}}=0.0095}$ Hrabec et al. (2016). We define quadratic and biquadratic interactions between magnetic layers using a custom exchange field and energy density, leading to the exchange energy in Eq. (1) (third and fourth terms on the right) that satisfies the form:
$\mathrm{\mathscr{E}_{\mathrm{exch}}=-\frac{1}{2}\mathrm{\overrightarrow{\bm{\mathrm{M}}}}\cdot\mathrm{\overrightarrow{\bm{\mathrm{H}}}}_{\mathrm{exch}}}.$ (5)
We can split the exchange field into quadratic and biquadratic terms as $\overrightarrow{\bm{H}}_{\mathrm{exch}}=\overrightarrow{\bm{H}}_{\mathrm{quad}}+\overrightarrow{\bm{H}}_{\mathrm{biquad}}$, which leads to the expressions for quadratic and biquadratic exchange fields given by:
$\mathrm{\overrightarrow{\bm{\mathrm{H}}}_{\mathrm{quad}}=-\frac{2J_{\mathrm{q}}}{M_{\mathrm{S1}}M_{\mathrm{S2}}}\tilde{\overrightarrow{\bm{\mathrm{M}}}}},$ (6a)
$\mathrm{\overrightarrow{\bm{\mathrm{H}}}_{\mathrm{biquad}}=-\frac{2J_{\mathrm{bq}}}{(M_{\mathrm{S1}}M_{\mathrm{S2}})^{2}}(\tilde{\overrightarrow{\bm{\mathrm{M}}}}\cdot\overrightarrow{\bm{\mathrm{M}}})\tilde{\overrightarrow{\bm{\mathrm{M}}}}}.$ (6b)
Here, $\tilde{\overrightarrow{\bm{M}}}$ is the swapped magnetization defined as $\tilde{\overrightarrow{\bm{M}}}(x,y,z,t)=\overrightarrow{\bm{M}}(x,y,-z,t)$.
We first performed magnetostatic simulations to generate the hysteresis loop. The simulation started with a static magnetic field of $+2000$ Oe along the long axis, swept down to $-2000$ Oe in steps of $20$ Oe, and ramped up back to $+2000$ Oe to complete the hysteresis loop. The built-in relaxation function in Mumax3 was used to find the equilibrium state of the system. The simulated hysteresis loops agree well with the results obtained by the VSM measurements, as shown in Fig. 1(c).
@@
In the following, the dynamics measurements are discussed and compared to micromagnetic simulations. The spin dynamics was measured for 11 samples with increasing thickness of Ru varying from 00 nm to $1$ nm by STFMR and thermal BLS techniques. Figure 2 compares the thermal BLS and STFMR spectra for samples with Ru thicknesses of 0.8 nm and 0.4 nm, both showing antiferromagnetic interaction in the magnetometry measurements (see Fig. S2). We notice two magnon modes in the spectra of all of our samples with negative RKKY interaction (as found in magnetometry measurements; see also SM). These two modes correspond to the optical and acoustic modes. Typically, these two modes are protected by symmetry, which means that they cross at the degeneracy point. As is evident from Fig. 2, that is not what we observe in the experiments: a clear avoided-level crossing is found at the degeneracy points, indicating a strong coupling between optical and acoustic modes. We observe the largest gap of 6 GHz for the sample with 0.4 nm of Ru thickness. The BLS spectrum qualitatively agrees with the STFMR spectrum for each sample. When comparing the BLS and STFMR measurements, we observe that the BLS intensity is less prominent for a particular mode, which corresponds to the optical mode, as discussed below. We speculate that this could be due to either the lower magnon population in the optical branch in the thermal BLS experiments or due to a lower BLS detection sensitivity to optical modes.
Micromagnetic simulations using Mumax3 reveal mode phase relations between the ferromagnetic layers in the dynamic spectra. Figure  3 simulates a synthetic ferrimagnetic structure with 0.8 nm Ru thickness, corresponding to experimental BLS and STFMR measurements in Figs. 2(a) and (c). The spectrum shows two hybridized magnon modes. Spatially-resolved micromagnetic modeling shows in-phase or out-of-phase modes away from the energy degeneracy point, corresponding to acoustic or optical modes [Fig. 3(b)]. At 40 Oe [marked ‘I’ in Fig. 3(a)], two modes are observed: a low-frequency mode at 2.4 GHz and a high-frequency mode at 8.3 GHz). The corresponding spatially-resolved simulations reveal the phase profile of the top and the bottom layer, see Fig. 3(b). As is evident from the figure, the top and the bottom layers are in-phase at 2.4 GHz (acoustic mode), while they are out-of-phase at 8.3 GHz (optical mode). The two modes are well-separated in frequency from one another; hence, they are uncoupled. At 1000 Oe [marked ‘III’ in Fig. 3(a)], the high-frequency mode is the acoustic mode, while the low-frequency mode is the optical mode, as can be seen from the spatially-resolved simulations shown in Fig. 3(b). Between 40 and 1000 Oe (field region between vertical dashed line ‘I’ and ‘III’), the low- and high-frequency modes exchange their character from acoustic to optical. As can be seen in Fig. 3(a), the two modes repel each other near the degeneracy point marked ‘II’, leading to an avoided level crossing and a gap in the spectrum (see SM for more details on the phase analysis). The corresponding spatially-resolved simulations reveal a ‘mixed’ phase relationship, Fig. 3(b). The modes are no longer purely acoustic or optical in nature. The opening of a gap at the degeneracy field is the signature of a strong coupling between the acoustic and optical modes. The results obtained by simulations (Fig. 3) support the experimental results (Fig. 2), corroborating a strong coupling between the modes. We argue that this strong coupling between acoustic and optical magnons is facilitated by exchange interaction induced by broken intrinsic symmetry in the studied synthetic ferrimagnet L D Landau and Lifshitz (1977).
The micromagnetic simulations also allow us to examine the dependence of gap size ($g$) and field ($H_{\mathrm{g}}$) on quadratic and biquadratic interaction: Figure 4(a) shows that $g$ and $H_{\mathrm{g}}$ linearly increase with the quadratic interaction constant $J_{\mathrm{q}}$. The corresponding simulated spectra are presented in the SM. Furthermore, we used the simulations to extract the dependence of $g$ and $H_{\mathrm{g}}$ on the biquadratic interaction $J_{\mathrm{bq}}$: Figure 4(b) demonstrates a minimal dependence of $g$ and $H_{\mathrm{g}}$ on $J_{\mathrm{bq}}$. The simulated $J_{\mathrm{b}}$ and $J_{\mathrm{bq}}$ corresponding to the experimental conditions found for a Ru thickness of 0.8 nm are shown as red data points in Figs. 4(a) and (b) for reference. In addition, we plot the experimentally obtained and simulated dependence between $g$ and $H_{\mathrm{g}}$ in Fig. 4(c). The overall trend matches remarkably well (please note that each experimental data point in Fig. 4(c) has different values for $J_{\mathrm{b}}$ and $J_{\mathrm{bq}}$, whereas different values of $J_{\mathrm{b}}$ were used for the simulated data points while $J_{\mathrm{bq}}$ was kept fixed). The micromagnetic simulations are particularly useful to determine $J_{\mathrm{q}}$ and $J_{\mathrm{bq}}$ even when $J_{\mathrm{q}}<2J_{\mathrm{bq}}$, in which case Eq. (3) is inapplicable due to incomplete antiferromagnetic alignment and undefined value of $H_{\mathrm{afm}}$. We discuss this approach in the following: If full antiferromagnetic alignment is experimentally achieved for a particular Ru thickness, $J_{\mathrm{q}}$ and $J_{\mathrm{bq}}$ can be obtained from the magnetometry results [Fig. 1 and Eqs. (3a) and (3b)]. At the same time, $H_{\mathrm{g}}$ can be determined from spin dynamics measurements such as STFMR or BLS. Since $H_{\mathrm{g}}$ is linearly proportional to $J_{\mathrm{q}}$, we can calculate the proportionality constant using linear regression for a given sample with a particular Ru thickness. This constant then allows us to determine $J_{\mathrm{q}}$ for any Ru thickness, even without full antiferromagnetic alignment. Finally, $J_{\mathrm{bq}}$ can be determined with the knowledge of $J_{\mathrm{q}}$ and Eq. (2a). This method is a powerful tool for determining the biquadratic exchange interaction of any system.

### v2

@@ III Results and Discussion
The relationship between the saturation magnetizations of the ferromagnetic layers can be established using the normalized magnetization of the forced ferromagnetic ($M_{\mathrm{ff}}$) and fully antiferromagnetic ($M_{\mathrm{afm}}$) states (the magnetizations at configuration ‘1’ and configuration ‘3’ in Fig. 1(d) respectively). The ratio of the saturation magnetizations can be expressed as follows:
$M_{\mathrm{r}}=\mathrm{\frac{M_{S2}}{M_{S1}}=\frac{M_{ff}-M_{afm}}{M_{ff}+M_{afm}}}.$ (4)
For a Ru thickness of $0.8\,\mathrm{nm}$, we extracted the critical fields from the VSM results [Fig. 1(d)] as $H_{0,\mathrm{ff}}=55\pm 1.5\,\mathrm{mT}$ and $H_{0,\mathrm{afm}}=13\pm 1\,\mathrm{mT}$, with corresponding normalized magnetizations $M_{\mathrm{ff}}=1.0$ and $M_{\mathrm{afm}}=0.302\pm 0.028$. From these values, the ratio of the saturation magnetization is calculated as $M_{\mathrm{r}}=0.536\pm 0.033$.
@@
Figures 2(a,b) show the STFMR spectra for samples with Ru thicknesses of 0.8 nm and 0.4 nm, both exhibiting antiferromagnetic exchange interaction as confirmed by magnetometry (see Fig. S2). We notice two magnon modes in the spectra of all of our samples with negative RKKY interaction (as found in magnetometry measurements; see also SM). These two modes correspond to the optical and acoustic modes (more details on the phase of the modes can be found below). Typically, these two modes are protected by symmetry, which means that they cross at the degeneracy point. As is evident from Fig. 2, that is not what we observe in the experiments: a clear avoided-level crossing is observed at the degeneracy points, indicating strong coupling between the optical and acoustic modes. The coupling gap is determined by fitting the resonance peaks in the STFMR spectra with standard Lorentzian functions (see SM for details) and identifying the minimum frequency separation between the resonances at a fixed magnetic field. The largest gap of 3.9 GHz is observed for the sample with a Ru thickness of 0.4 nm. The enhanced signal amplitude at 3.8 GHz is a measurement artifact and does not carry physical significance.
We model the magnetization dynamics by treating the two ferromagnetic layers as coupled macrospins, whose time evolution is governed by the Landau–Lifshitz–Gilbert (LLG) equation Landau and Lifshitz (1935); Gilbert (2004); Stancil and Prabhakar (2008):
$\frac{d\mathbf{m}_{\mathrm{1}}}{dt}$ $=-\mu_{0}\gamma\mathbf{m}_{\mathrm{1}}\times\Big[H_{0}\hat{\mathbf{y}}-H_{\mathrm{E1}}\mathbf{m}_{\mathrm{2}}$
$\qquad-H_{\mathrm{B1}}(\mathbf{m}_{\mathrm{1}}\cdot\mathbf{m}_{\mathrm{2}})\mathbf{m}_{\mathrm{2}}-M_{\mathrm{S1}}(\mathbf{m}_{\mathrm{1}}\cdot\hat{\mathbf{z}})\hat{\mathbf{z}}\Big$]. (5a)
$\frac{d\mathbf{m}_{\mathrm{2}}}{dt}$ $=-\mu_{0}\gamma\mathbf{m}_{\mathrm{2}}\times\Big[H_{0}\hat{\mathbf{y}}-H_{\mathrm{E2}}\mathbf{m}_{\mathrm{1}}$
$\qquad-H_{\mathrm{B2}}(\mathbf{m}_{\mathrm{1}}\cdot\mathbf{m}_{\mathrm{2}})\mathbf{m}_{\mathrm{1}}-M_{\mathrm{S2}}(\mathbf{m}_{\mathrm{2}}\cdot\hat{\mathbf{z}})\hat{\mathbf{z}}\Big$]. (5b)
Here, $\mathbf{m}_{1}$ and $\mathbf{m}_{2}$ denote the macrospin unit vectors of the ferromagnetic layers, with their dynamics described by
$\frac{d\mathbf{m}_{\mathrm{i}}}{dt}=i\omega\,\delta\mathbf{m}_{\mathrm{i}},$
where $\delta\mathbf{m}_{\mathrm{i}}$ is the unit vector of the dynamic deviation of the magnetization unit vector. The effective fields $H_{\mathrm{E\,i}}$ and $H_{\mathrm{B\,i}}$ due to quadratic and biquadratic exchange are defined as
$H_{\mathrm{E\,i}}=\frac{J_{\mathrm{q}}}{M_{\mathrm{S\,i}}},\qquad H_{\mathrm{B\,i}}=\frac{2J_{\mathrm{bq}}}{M_{\mathrm{S\,i}}}.$
Finally, the torque terms proportional to the saturation magnetizations $M_{\mathrm{S1}}$ and $M_{\mathrm{S2}}$ in Eqs. 5a and 5b parameterize the shape-dependent demagnetizing fields of the NiFe and CoFe layers, respectively.
After some algebra (see SM), the nontrivial solutions are obtained by requiring that the determinant of the coefficient matrix vanishes:
$\begin{vmatrix}i\omega&\mu_{0}\gamma(M_{1}+H_{\mathrm{1,eq}})&0&\frac{\mu_{0}\gamma}{M_{1}}(J_{\mathrm{q}}+2J_{\mathrm{bq}}\cos{\alpha})\\ -\mu_{0}\gamma\left(\frac{2J_{\mathrm{bq}}}{M_{1}}\sin^{2}{\alpha}+H_{\mathrm{1,eq}}\right)&i\omega&\frac{\mu_{0}\gamma}{M_{1}}\left[2J_{\mathrm{bq}}(\sin^{2}{\alpha}-\cos^{2}{\alpha})-J_{\mathrm{q}}\cos{\alpha}\right]&0\\ 0&\frac{\mu_{0}\gamma}{M_{2}}(J_{\mathrm{q}}+2J_{\mathrm{bq}}\cos{\alpha})&i\omega&\mu_{0}\gamma(M_{2}+H_{\mathrm{2,eq}})\\ \frac{\mu_{0}\gamma}{M_{2}}\left[2J_{\mathrm{bq}}(\sin^{2}{\alpha}-\cos^{2}{\alpha})-J_{\mathrm{q}}\cos{\alpha}\right]&0&-\mu_{0}\gamma\left(\frac{2J_{\mathrm{bq}}}{M_{2}}\sin^{2}{\alpha}+H_{\mathrm{2,eq}}\right)&i\omega\\ \end{vmatrix}=0,$
where $H_{\mathrm{i,eq}}$ denotes the effective magnetic field experienced by layer $i$, and $\alpha=\phi_{\mathrm{a}}+\phi_{\mathrm{b}}$ is the equilibrium interlayer angle.
Numerical solutions of the determinant yield the magnon branches for a given set of parameters, and the experimental spectra can be fitted within this framework to extract the corresponding material parameters. Fitting the STFMR spectra with the macrospin model is complicated by the hidden variables $\phi_{\mathrm{a}}$ and $\phi_{\mathrm{b}}$, which describe the equilibrium magnetization orientations of the two ferromagnetic layers and cannot be measured directly. Experimentally, we obtain the resonance frequency as a function of the external magnetic field, while the angles $\phi_{\mathrm{a}}$ and $\phi_{\mathrm{b}}$ must be determined self-consistently from the model. The fitting algorithm and corresponding Python code are provided in the SM.
For a Ru thickness of $0.8\,\mathrm{nm}$, the fitted parameters are $M^{\mathrm{fit}}_{\mathrm{S1}}=1.82\times 10^{6}\penalty 10000\ \mathrm{A/m}$, $M^{\mathrm{fit}}_{\mathrm{r}}=0.527$, $J^{\mathrm{fit}}_{\mathrm{q}}=29.8\penalty 10000\ \mathrm{kJ/m^{3}}$, and $J^{\mathrm{fit}}_{\mathrm{bq}}=2.16\penalty 10000\ \mathrm{kJ/m^{3}}$. The fitted macrospin magnon spectrum is overlaid with the STFMR data in Fig. 2(a). An analogous procedure was followed for the sample with a Ru thickness of $0.4\penalty 10000\ \mathrm{nm}$, shown in Fig. 2(b), demonstrating good agreement in both cases. Using the material parameters obtained from the dynamic macrospin model, we compare the calculated hysteresis loop with the experimental VSM data shown in Fig. 1(d) and find good overall agreement. Finally, these parameters also yield the knee fields from Eqs. (2a) and (2b), resulting in $H^{\mathrm{fit}}_{0,\mathrm{ff}}=54.4\penalty 10000\ \mathrm{mT}$ and $H^{\mathrm{fit}}_{0,\mathrm{afm}}=12.6\penalty 10000\ \mathrm{mT}$, which are consistent with the values extracted from the VSM measurements ($H_{0,\mathrm{ff}}=55\pm 1.5\,\mathrm{mT}$ and $H_{0,\mathrm{afm}}=13\pm 1\,\mathrm{mT}$) within experimental uncertainty.
To gain a better understanding of the experimental results, we performed micromagnetic modeling using the graphics processor unit (GPU)-accelerated program MuMax3 Vansteenkiste et al. (2014); Exl et al. (2014). Our structure was modeled as $128\times 64\times 3$ cells with an individual cell size of $4.0$ nm × $\times\penalty 10000\ 4.0$ nm × $\times\penalty 10000\ 4.0$ nm, with the middle layer representing the spacer. Each layer was defined into a region to set material parameters separately. The values for the exchange stiffness of CoFe and NiFe used in the simulation were $3.0\times 10^{-11}$ J/m and $1.3\times 10^{-11}$ J/m, respectively Parreiras and Martins (2015). This corresponds to exchange lengths for NiFe and CoFe of 5.29 nm and 4.32 nm, respectively Parreiras and Martins (2015). Crystalline anisotropy was considered negligible. The following Gilbert damping constants were used for CoFe and NiFe, respectively: $\mathrm{\alpha_{\mathrm{CoFe}}=0.005}$ Weber et al. (2019) and $\mathrm{\alpha_{\mathrm{NiFe}}=0.0095}$ Hrabec et al. (2016). We define the quadratic and biquadratic interlayer exchange interactions using custom expressions for the exchange field and the corresponding energy density. These expressions are constructed to reproduce the exchange energy given in Eq. (1) (the third and fourth terms on the right-hand side) and to ensure that the resulting magnetization dynamics are consistent with the LLG equation. We define the quadratic and biquadratic exchange fields as (see SM):
$\mathrm{\boldsymbol{\mathrm{H}}_{\mathrm{quad}}=-\frac{J_{\mathrm{q}}}{M_{\mathrm{S1}}M_{\mathrm{S2}}}\widetilde{\boldsymbol{\mathrm{M}}}}$, (7a)
$\mathrm{\boldsymbol{\mathrm{H}}_{\mathrm{biquad}}=-\frac{2J_{\mathrm{bq}}}{(M_{\mathrm{S1}}M_{\mathrm{S2}})^{2}}(\widetilde{\boldsymbol{\mathrm{M}}}\cdot\boldsymbol{\mathrm{M}})\widetilde{\boldsymbol{\mathrm{M}}}}$. (7b)
and corresponding energy density terms as:
$\mathrm{\mathscr{E}_{\mathrm{quad}}=-\mathrm{\boldsymbol{\mathrm{M}}}}\cdot\mathrm{\boldsymbol{\mathrm{H}}}_{\mathrm{quad}}.$ (8a)
$\mathrm{\mathscr{E}_{\mathrm{biquad}}=-\frac{1}{2}\mathrm{\boldsymbol{\mathrm{M}}}}\cdot\mathrm{\boldsymbol{\mathrm{H}}}_{\mathrm{biquad}}.$ (8b)
Here, $\widetilde{\boldsymbol{M}}$ is the swapped magnetization defined as $\widetilde{\boldsymbol{M}}(x,y,z,t)=\boldsymbol{M}(x,y,-z,t)$. For the simulations, the saturation magnetizations and exchange coupling constants were taken from macrospin model fits to the STFMR spectra.
We first performed magnetostatic simulations to generate the hysteresis loop. We started the simulation with a static magnetic field of +200+200 mT along the long axis, swept down to 200-200 mT in steps of $2$ mT, and ramped up back to +200+200 mT to complete the hysteresis loop. The built-in relaxation function in MuMax3 was used to find the equilibrium state of the system. The simulated hysteresis loops show good agreement with both the VSM measurements and the macrospin model, as illustrated in Fig. 1(d) for a Ru thickness of $0.8\,\mathrm{nm}$. MuMax3 simulations of the equilibrium angles of the ferromagnetic layers with respect to the external field, $\phi_{\mathrm{a}}$ (CoFe) and $\phi_{\mathrm{b}}$ (NiFe), are plotted in Fig. 1(e) as a function of the applied field during field ramp-down. Note that -180° and 180° correspond to the same magnetization direction. In some regions, the blue data points (magnetization direction of CoFe) are obscured by the overlapping orange data points (magnetization direction of NiFe). This plot provides direct insight into the transitions observed in the magnetostatic measurements and calculations shown in Fig. 1(d). In particular, the transition from configuration ‘1’ to configuration ‘2’ corresponds to a spin-flop transition, whereas the transition from configuration ‘3’ to configuration ‘4’ corresponds to a spin-flip transition.
Figure 4: (a) Coupling gap $g$ (blue) and degeneracy field $H_{\mathrm{g}}$ (green) obtained from macrospin calculations (solid lines) and micromagnetic simulations (dashed lines) as a function of the quadratic exchange constant $J_{\mathrm{q}}$ at a fixed biquadratic exchange constant of $J_{\mathrm{bq}}=2.16\penalty 10000\ \mathrm{kJ/m^{3}}$. The degeneracy field $H_{\mathrm{g}}$ is defined as the magnetic field at which the frequency separation between the high- and low-frequency modes is minimized, yielding the coupling gap $g$. The uncertainty (error bar) in the simulated $H_{g}$ arises from the finite field resolution, given by the field step size (2 mT in this case). The uncertainty in the simulated $g$ is determined by the finite frequency resolution of the FFT, set by the simulation runtime (100 MHz for 10 ns), and is multiplied by $\sqrt{2}$ to account for error propagation. (b) Same as (a), but shown as a function of the biquadratic exchange constant $J_{\mathrm{bq}}$ at a fixed quadratic exchange constant of $J_{\mathrm{q}}=29.8\penalty 10000\ \mathrm{kJ/m^{3}}$. (c–d) The b coefficient in the sublinear power-law dependence on $J_{\mathrm{bq}}$ and $J_{\mathrm{q}}$, respectively.
Dynamic simulations using MuMax3 reveal the phase relations between the ferromagnetic layers in the dynamic spectra. Figures 2(c,d) show simulated spectra for Ru thicknesses of 0.8 nm and 0.4 nm, corresponding to the experimental STFMR measurements in Figs. 2(a,b). We started the simulation with a static magnetic field of +200 mT and swept down to 0 mT in steps of 1 mT, with the initial magnetization state at each field taken from magnetostatic simulations. A sinc pulse with an amplitude of $0.1\penalty 10000\ \mathrm{mT}$ and a cutoff frequency of 25 GHz was applied along all three spatial axes to simultaneously excite both the acoustic and optical modes Subedi et al. (2025), while a static external field was applied along the long axis of the sample. The RF field direction and phase were chosen to minimize numerical artifacts. At each field step, the dynamics were evolved for 10 ns, and the total magnetization of the coupled layers was Fourier transformed to obtain the spectra. Macrospin model calculations are overlaid on the simulations and show excellent agreement. The spectra show two hybridized magnon branches similar to the experimentally obtained spectra, one branch appearing more prominent than the other, consistent with the general observation that the acoustic mode is readily excited whereas the optical mode is more difficult to excite, particularly in the saturated regime Xiao et al. (2022). Importantly, the RF field direction and phase do not affect the spectra (see SM).
We further analyzed the FFT phase to determine the magnon mode relations at different fields. At 4 mT [configuration ‘3’ in Fig. 1(d,e); see Fig. 3(b)], when the layers are antiferromagnetically aligned, the FFT magnitude (top panel) exhibits two peaks at 1.16 GHz and 7.68 GHz in both ferromagnetic layers, marked by vertical lines. At the lower frequency branch (1.16 GHz), the CoFe and NiFe phases (bottom panel) are $151.6^{\circ}$ and $151.3^{\circ}$, marked by solid horizontal lines, differing by only $0.3^{\circ}$, consistent with an in-phase acoustic mode. At the high-frequency branch (7.68 GHz), the phases are $14.4^{\circ}$ (CoFe) and $157.3^{\circ}$ (NiFe), marked by dashed horizontal lines, differing by $171.7^{\circ}$, characteristic of an out-of-phase optical mode.
At the transition field $H_{\mathrm{0,afm}}=12\penalty 10000\ \mathrm{mT}$ [transition point from configuration ‘2’ to configuration ‘3’ in Fig. 1(d,e); Fig. 3(c)], the phase differences remain close to $1^{\circ}$ (acoustic) and $189^{\circ}$ (optical). Near the degeneracy point, the FFT phase exhibits a rapid variation. For this reason, we performed simulations with higher magnetic-field resolution, using a step size of 0.2 mT. At 37.4 mT [Fig. 3(d)], we find an avoided level crossing and a gap in the spectrum, where the phase differences increase to $40^{\circ}$ and $230^{\circ}$, respectively, neither close to what we would expect from an acoustic or optical mode. The modes are no longer purely acoustic or optical in nature, indicating hybridization of the modes. At the second transition field $H_{\mathrm{0,ff}}=52\penalty 10000\ \mathrm{mT}$ [transition point from configuration ‘1’ to configuration ‘2’ in Fig. 1(d,e); Fig. 3(e)], the phase differences switch to $188^{\circ}$ (low frequency) and $6^{\circ}$ (high frequency), corresponding to an interchange of acoustic and optical character of the two modes. In the saturated regime [configuration ‘1’; Fig. 3(f)], at 100 mT the phase differences are $197^{\circ}$ and $7^{\circ}$, again indicating optical and acoustic modes, respectively. Finally, spatially resolved simulations at 12 mT [Fig. 3(g)] show a uniform phase within each layer, with no internal phase variation.
The opening of a gap at the degeneracy field is the signature of a strong coupling between the acoustic and optical modes. The results obtained by simulations support the experimental results (Fig. 2), corroborating a strong coupling between the modes. We argue that this strong coupling between acoustic and optical magnons is facilitated by exchange interaction induced by broken intrinsic symmetry in the studied synthetic ferrimagnet L D Landau and Lifshitz (1977).
The macrospin model and micromagnetic simulations further reveal the dependence of the coupling gap $g$ and the corresponding field $H_{\mathrm{g}}$ on the quadratic and biquadratic exchange interactions. As shown in Figs. 4(a,b), $H_{\mathrm{g}}$ increases linearly with both $J_{\mathrm{q}}$ and $J_{\mathrm{bq}}$. In contrast, the gap size $g$ exhibits a sublinear power-law dependence on $J_{\mathrm{q}}$, with the scaling behavior governed by the strength of $J_{\mathrm{bq}}$.
The dependence of $g$ on $J_{\mathrm{q}}$ can be modeled as:
$g=a\,J_{\mathrm{q}}^{\,b}+c,$ (9)
where the fitting parameters $a$, $b$, and $c$ depend on the biquadratic exchange interaction. For $J_{\mathrm{bq}}=2.16\penalty 10000\ \mathrm{kJ/m^{3}}$, we obtain a fitted exponent of $b=0.79$ [Fig. 4(a)]. More generally, the exponent $b$ increases with increasing $J_{\mathrm{bq}}$, starting from $b=0.5$ in the absence of biquadratic exchange ($J_{\mathrm{bq}}=0$) and approaching a linear scaling regime ($b=1$) at large $J_{\mathrm{bq}}$, as shown in Fig. 4(c).
The dependence of the gap $g$ on the biquadratic exchange interaction $J_{\mathrm{bq}}$ is regime dependent. For $J_{\mathrm{q}}<2J_{\mathrm{bq}}$, where a fully antiferromagnetic alignment is not energetically accessible, $g$ follows a sublinear power-law dependence on $J_{\mathrm{bq}}$, with the scaling exponent determined by $J_{\mathrm{q}}$. In contrast, for $J_{\mathrm{q}}>2J_{\mathrm{bq}}$, where a fully antiferromagnetic state becomes accessible [see Eq. (2b)], the gap $g$ decreases with increasing $J_{\mathrm{bq}}$ and deviates from the sublinear power-law behavior.
For $J_{\mathrm{q}}=30\penalty 10000\ \mathrm{kJ/m^{3}}$, we obtain an approximately linear dependence of $g$ on $J_{\mathrm{bq}}$ in the regime $J_{\mathrm{q}}>2J_{\mathrm{bq}}$, with a fitted exponent of $b=1.02$ [Fig. 4(b)]. More generally, the exponent $b$ increases with increasing $J_{\mathrm{q}}$, starting from $b=0.5$ in the absence of quadratic exchange ($J_{\mathrm{q}}=0$) and approaching a linear scaling regime ($b=1$) at large $J_{\mathrm{q}}$, as summarized in Fig. 4(d). Further details are provided in the SM.